Circular No. 5753

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 Y. Tanaka (Institute of Space and Astronautical Science) and the

ASCA team communicate: “The x-ray astronomy satellite ASCA (Advanced Satellite for Cosmology and Astrophysics), launched on Feb. 20, observed SN 1993J from Apr. 5.25 to Apr. 6.03 UT for an exposure time of ∼ 37 000 s with the gas scintillation imaging spectrometers and the x-ray CCD cameras. A bright x-ray source was found at the optically determined position of SN 1993J within an accuracy of 0′.1. More than 10 000 photons were collected from the source. For an assumed distance of 3.2 Mpc, the source luminosity was ∼ 5 × 10^{39} erg/s in the range 1-10 keV. The energy spectrum is hard; if it is a thermal bremsstrahlung spectrum, kT > 10 keV and N_{H} ∼ 10^{21} cm^{−2}. More observations will follow.”

PKS 0528+134 The Compton Observatory EGRET Team (P. Sreekumar, B. L. Din-

gus, D. L. Bertsch, C. E. Fichtel, J. A. Esposito, R. C. Hartman, S. D. Hunter, J. R. Mattox, and D. J. Thompson, Goddard Space Flight Center, NASA; D. A. Kniffen, Hampden-Sydney College; J. Chiang, J. Fierro, Y. C. Lin, P. F. Michelson, and P. L. Nolan, Stanford University; E. Schneid, Grumman Aerospace Corporation; and K. Brazier, G. Kanbach, H. A. Mayer-Hasselwander, C. von Montigny, Max-Planck-Institut für Extraterrestrische Physik) communicates: “We report detection of γ-rays from the quasar PKS 0528+134 during the observation interval Mar. 23-29 at a level approximately three times the observed intensity during the early part of the GRO mission (1991 May 16-30; Hunter et al. 1993, to appear in Ap.J. 409). This makes PKS 0528+134 as bright as the Crab Pulsar in high γrays. Additional EGRET observations of this region in the sky are planned during 1993 May 13-25.”

CORRIGENDA On IAUC 5728, ‘GRO J0422+32’, line 10, for (equivalent width 2 nm)

read (equivalent width 0.2 nm)

On IAUC 5733, SN 1993J, line 3, for Mar. 30 read Mar. 29.876 UT On IAUC 5750, line 10, for NH column density read N_{H} column density On IAUC 5752, SN 1993J, last line, for 12 mJy at 20 cm.” read 12 mJy

at 2 cm.”

1993 April 7 Daniel W. E. Green